a rental car costs $25 plus a fixed charge per miles driven. The total charge for 320 miles of use was $89. write an equation for the cost, C (in dollars) in terms of the miles driven, x.
Answer by mananth(16946)

a rental car costs $25 plus a fixed charge per miles driven. The total charge for 320 miles of use was $89. write an equation for the cost, C (in dollars) in terms of the miles driven, x.
..
fixed charges = $25
...
89 -25 = $ 64 charge for 320 miles driven
..
320 mi 64
1 mile = x
x= 64/320
x= $0.2 / mile
..
the Cost C = 25 + 0.2x
...
m.ananth@hotmail.ca